# Slender Trees and the Approximation Property

### v1

@@ 4. Ineffability Witnesses
    $\mathbb{P}$ is of size $\leq\lambda$ and $\kappa$-cc.,
@@
      $[M\cap\lambda]^{<M\cap\kappa}\subseteq M$.
@@
      Whenever $G$ is $\mathbb{P}$-generic over $V$, $\pi[G\cap M]$ is $\pi(\mathbb{P})$-generic over $V$ and the pair $(V[\pi[G\cap M]],V[G])$ has the ${<}\,\pi(\delta)$-approximation property.
@@
[…]
@@
    If there exists $M\prec H(\Theta)$ (with Mostowski-Collapse $\pi$) such that $M\cap\kappa\in\kappa$, $M\cap\lambda=a$, $[M\cap\lambda]^{<M\cap\kappa}\subseteq M$, $\kappa,\lambda,\dot{f},\dot{F}\in M$ and for some $\pi(\mathbb{P})$-name $\dot{x}_{a}$ for a subset of $\pi[M\cap\lambda]$ and a condition $p_{a}\leq p$, $p_{a}\Vdash\dot{f}(M\cap\lambda)=\pi^{-1}[\dot{x}_{a}^{\pi[\Gamma\cap M]}]$, let
@@
Let $M\in[H(\Theta)]^{<\kappa}$ be a $\lambda$-ineffability witness for $\kappa$ with respect to $e$ as in the requirements. Let $\pi\colon M\to N$ be the Mostowski-Collapse and $a:=M\cap\lambda$ as well as $\theta:=\pi[a]$. We will show that case (1) holds. Let $G_{0}$ be $\mathbb{P}$-generic containing $p$ and $G_{0}^{\prime}:=\pi[G_{0}\cap M]$. By Lemma 4.6, $\pi$ extends to $\pi\colon M[G_{0}]\to N[G_{0}^{\prime}]$ and $[\theta]^{<\nu}\cap V[G_{0}^{\prime}]\subseteq N[G_{0}^{\prime}]$. Furthermore, $M[G_{0}]\prec H(\Theta)^{V[G_{0}]}$, so $M[G_{0}]\cap H(\Theta^{\prime})$ is closed under $\dot{F}^{G_{0}}$ and therefore witnesses the slenderness of $\dot{f}^{G_{0}}$. Assume $\pi[\dot{f}^{G_{0}}(a)]\notin V[G_{0}^{\prime}]$. By the ${<}\,\pi(\delta)$-approximation property there is $z\in V[G_{0}^{\prime}]$ with ordertype ${<}\,\pi(\delta)$ such that $\pi[\dot{f}^{G_{0}}(a)]\cap z\notin V[G_{0}^{\prime}]$. We can assume $z\subseteq a\pi[a]$, so we have $z\in N[G_{0}^{\prime}]$ and $\pi^{-1}(z)=\pi^{-1}[z]\in M[G_{0}]$. Because $M[G_{0}]\cap H(\Theta^{\prime})$ witnesses the slenderness of $\dot{f}^{G_{0}}$, we have
$\pi[\dot{f}^{G_{0}}(a)]\cap z=\pi[\dot{f}^{G_{0}}(a)]\cap\pi[\pi^{-1}[z]]=\pi[\underbrace{\dot{f}^{G_{0}}(a)\cap\pi^{-1}[z]}_{\in M[G_{0}],\;\subseteq M[G_{0}]}]=\pi(\dot{f}^{G_{0}}(a)\cap\pi^{-1}[z])\in N[G_{0}^{\prime}]$
a contradiction. So $\pi[\dot{f}^{G_{0}}(a)]\in V[G_{0}^{\prime}]$. Ergo there is $p_{a}\leq p$ as well as a $\pi(\mathbb{P})$-name $\dot{x}_{a}$ for a subset of $\theta$ such that $p_{a}\Vdash\pi[\dot{f}(a)]=\dot{x}_{a}^{\pi[\Gamma\cap M]}$ which is what we wanted to show. Now our aim is to show that $p_{a}$ forces $M[\Gamma]$ to be a $\lambda$-ineffability witness for $\kappa$ with respect to $\dot{f}$. So let $G_{1}$ be a $\mathbb{P}$-generic filter containing $p_{a}$. It is clear that $\kappa,\lambda,\dot{f}^{G_{1}}\in M[G_{1}]$. By assumption there is $b_{e}\in M$ such that $b_{e}\cap M=e(M\cap\lambda)$. Define
@@
all that is left is to show $b_{f}\cap M=\dot{f}^{G_{1}}(M\cap\lambda)=\dot{f}^{G_{1}}(M[G_{1}]\cap\lambda)$ (where the last equality holds by Lemma 4.6). Let $\beta\in b_{f}\cap M$. By elementarity there is $\alpha\in G_{1}\cap M$ such that $\langle\alpha,\beta\rangle\in b_{e}$ and we have $\langle\alpha,\beta\rangle\in b_{e}\cap M=e(a)=\pi^{-1}[g(a)]$. So $\pi(\langle\alpha,\beta\rangle)=\langle\pi(\alpha),\pi(\beta)\rangle\in g(a)$. By the definition, $\pi(\alpha)\Vdash\pi(\check{\beta})\in\dot{x}_{a}$. Hence $\pi(\beta)\in\dot{x}_{a}^{\pi[G_{1}\cap M]}$ and $\beta\in\dot{f}^{G_{1}}(a)$. Let $\beta\in\dot{f}^{G_{1}}(M\cap\lambda)$, so $\pi(\beta)\in\dot{x}_{a}^{\pi[G_{1}\cap M]}$. Thus there exists $\pi(\alpha)\in\pi[G_{1}\cap M]$ such that $\pi(\alpha)\Vdash\pi(\beta)\in\dot{x}_{a}$. Hence $\pi(\langle\alpha,\beta\rangle)\in g(a)$ and $\langle\alpha,\beta\rangle\in e(a)=b_{e}\cap M$. Ergo $\beta\in b_{f}\cap M$. ∎
@@
Let $\mathbb{P}:=\operatorname{Add}(\tau,\kappa)$. Let $e$ be any $(\kappa,\lambda)$-list, $\Theta$ large and $x\in[H(\Theta)]^{<\kappa}$. Find a $\lambda$-ineffability witness $M$ for $\kappa$ as in Lemma 4.3 with $\tau^{+}+1\subseteq M$. It follows that $\pi(\mathbb{P})=\operatorname{Add}(\tau,M\cap\kappa)$. If $G$ is $\mathbb{P}$-generic, $\pi[G\cap M]$ is $\pi(\mathbb{P})$-generic and $V[G]$ is an extension of $V[\pi[G\cap M]]$ by $\operatorname{Add}(\tau,\kappa\smallsetminus(M\cap\kappa))$. In $V[\pi[G\cap M]]$, this forcing is $\tau^{+}$-Knaster and thus has the ${<}\,\tau^{+}=\pi(\tau^{+})$-approximation property (viewing it as an ordering on the product ×$\mathbb{P}\times\{0\}$ which is clearly iteration-like). ∎
@@ 5. Controlling the Slenderness at $\kappa$
We will start with the easiest construction regarding $\operatorname{ISP}$: In this section we will define a forcing $\mathbb{M}_{0}(\tau,\mu,\kappa)$ which forces $2^{\tau}=\mu^{+}=\kappa$ and $\operatorname{ISP}(\tau^{+},\kappa,\lambda)$ (if $\kappa$ is $\lambda$-ineffable), answering Problem 4.31 given by Mohammadpour in [11]. Additionally, it will force $\neg\operatorname{ISP}(\tau,\kappa,\kappa)$. Thus we are able to control exactly how slender lists must be to have guaranteed ineffable branches. For this section, fix regular cardinals $\tau<\mu<\kappa$ such that $\tau^{<\tau}=\tau$ and $\kappa$ is inaccessible. For technical reasons, for any ordinal $\gamma$, we let $\operatorname{Add}(\tau,\gamma)$ be the ${<}\,\tau$-supported product of $\operatorname{Add}(\tau)$ of length $\gamma$, i.e. conditions are partial functions $p$ on $\gamma$ of size ${<}\,\tau$ such that for all $\alpha\in\operatorname{dom}(p)$, $p(\alpha)\in\operatorname{Add}(\tau)$, ordered in the usual way.
@@
    $q$ is a ${<}\,\mu$-sized partial function on $\nu$ such that whenever $\gamma\in\operatorname{dom}(q)$, $\gamma=\delta+1$ for an inaccessible cardinal $\delta$ and $q(\gamma)$ is an $\operatorname{Add}(\tau,\gamma)$-name for a condition in $\dot{\operatorname{Coll}}(\check{\mu},\check{\delta})$.
@@
The forcing $\mathbb{M}_{0}$ was used by Levine in [9] to give an easier construction of a model with a disjoint stationary sequence on an arbitrary double successor cardinal and disjoint stationary sequences on two successive cardinals. We can obviously view $\mathbb{M}_{0}(\tau,\mu,\nu)$ as an ordering on some product $\operatorname{Add}(\tau,\nu)\times\mathbb{T}(\tau,\mu,\nu)$, which is based.
@@
    $\mathbb{M}_{0}(\tau,\mu,\nu)$ is ${<}\,\tau$-directed closed.
@@
    The term ordering on $\mathbb{M}_{0}(\tau,\mu,\nu)$ is ${<}\,\mu$-closed,
@@
The statements (1) and (2) are standard. We give the proofs for (3) and (4) in order to familiarize ourselves with the definitions. Regarding the base ordering, we observe that $(p,q)\leq(p^{\prime},q)$ if and only if $p\leq p^{\prime}$ in the ordering on $\operatorname{Add}(\tau,\gamma)$. So the base ordering is equal to $\operatorname{Add}(\tau,\gamma)$ which is $(2^{<\tau})^{+}=\tau^{+}$-Knaster. To show (4), let $(p,q_{\alpha})_{\alpha<\delta}$ (for $\delta<\mu$) be a descending sequence. Let $x:=\bigcup_{\alpha<\delta}\operatorname{dom}(q_{\alpha})$. $|x|<\mu$. On $x$, define a function $q$ as follows: Given $\gamma\in x$, let $\alpha_{0}$ be such that $\gamma\in\operatorname{dom}(q_{\alpha_{0}})$ (so $\gamma\in\operatorname{dom}(q_{\alpha})$ for all $\alpha\geq\alpha_{0})$. Then $(q_{\alpha}(\gamma))_{\alpha_{0}\leq\alpha<\delta}$ is forced by $p$ to be a descending sequence in $\dot{\operatorname{Coll}}(\check{\mu},\check{\delta})$, so by the maximum principle we can fix $q(\gamma)$ which is forced to be a lower bound. It follows that $(p,q)$ is a lower bound of $(p,q_{\alpha})_{\alpha<\delta}$. Regarding the projection property, given $(p^{\prime},q^{\prime})\leq(p,q)$, let $q^{\prime\prime}$ be a function with domain $\operatorname{dom}(q^{\prime})$ such that for all $\alpha\in\operatorname{dom}(q^{\prime})$, $q^{\prime\prime}(\alpha)$ is forced by $p^{\prime}$ to be equal to $q^{\prime}(\alpha)$ and by conditions incompatible to $p^{\prime}$ to be equal to $q(\alpha)$ (or $\emptyset$, if $q(\alpha)$ is not defined). Clearly, $q^{\prime\prime}$ is as required. $\mathbb{M}_{0}$ has the refinement property since stronger conditions force more. For the mixing property, let $(p,q_{0}),(p,q_{1})\leq(p,q)$. Let $p_{0},p_{1}$ be extensions of $p$ such that $p_{0}(0)$ and $p_{0}(1)$ are incompatible. Construct a function $q^{\prime}$ with domain $x:=\operatorname{dom}(q_{0})\cup\operatorname{dom}(q_{1})$ As follows: Let $\alpha\in x$. Then $p_{0}\upharpoonright\alpha$ and $p_{1}\upharpoonright\alpha$ are incompatible since $\alpha$ is the successor of a cardinal. So we can let $q^{\prime}(\alpha)$ be such that it is forced by $p_{0}\upharpoonright\alpha$ to be equal to $q_{0}(\alpha)$ (or $\emptyset$, if $q_{0}(\alpha)$ is not defined) and by conditions incompatible with $p_{0}\upharpoonright\alpha$ (in particular, $p_{1}\upharpoonright\alpha$) to be equal to $q_{1}(\alpha)$ (or $\emptyset$, if $q_{1}(\alpha)$ is not defined). It follows that $(p,q^{\prime})\leq(p,q)$ and for $i=0,1$, $(p_{i},q^{\prime})\leq(p,q_{i})$. ∎
@@
    $q$ is a ${<}\,\mu$-sized partial function on $\kappa\smallsetminus\nu$ such that for all $\gamma\in\operatorname{dom}(q)$, $\gamma=\delta+1$ for an inaccessible cardinal $\nu$ and $q(\gamma)$ is an $\operatorname{Add}(\tau,\gamma\smallsetminus\nu)$-name for a condition in $\dot{\operatorname{Coll}}(\check{\mu},\check{\nu})$.
@@
    $\mathbb{M}_{0}(G,\tau,\mu,\kappa\smallsetminus\nu)$ is ${<}\,\tau$-directed closed.
@@
    The term ordering on $\mathbb{M}_{0}(G,\tau,\mu,\kappa\smallsetminus\nu)$ is ${<}\,\mu$-closed,
@@
Let $\nu\in(\mu,\kappa)$ be inaccessible and let $G$ be $\mathbb{M}_{0}(\tau,\mu,\nu)$-generic. In $V[G]$, the forcing $\mathbb{M}_{0}(G,\tau,\mu,\kappa\smallsetminus\nu)$ has the ${<}\,\tau^{+}$-approximation property.
@@
As is common, we define $\iota((p,q)):=((p\upharpoonright\nu,q\upharpoonright\nu),(\check{p\upharpoonright[\nu,\kappa)},\overline{q}))$, where $\overline{q}$ is $\mathbb{M}_{0}(\tau,\mu,\nu)$-name for a function on $\kappa\smallsetminus\nu$ given by recomputing $q\upharpoonright[\nu,\kappa)$. It is clear that $\iota$ preserves $\leq$. Let $((p_{0},q_{0}),\sigma)$ be a condition in $\mathbb{M}_{0}(\tau,\mu,\nu)*\mathbb{M}_{0}(\Gamma,\tau,\mu,\kappa\smallsetminus\nu)$. By strengthening $(p_{0},q_{0})$ if necessary, we can assume that it forces $\sigma=(\check{p_{1}},\sigma^{\prime})$ for some $p_{1}\in\operatorname{Add}(\tau,\kappa\smallsetminus\nu)$ (because $\mathbb{M}_{0}(\tau,\mu,\nu)$ is ${<}\,\tau$-directed closed). $\sigma^{\prime}$ is forced to be a ${<}\,\mu$-sized sequence of ${<}\,\mu$-sized subsets of $V$, so we can (again, strengthening $(p_{0},q_{0})$ if necessary) assume that $\sigma^{\prime}$ is an $\operatorname{Add}(\tau,\nu)$-name. Lastly, by the $\mu$-cc. of $\operatorname{Add}(\tau,\nu)$, there is a set $x\in V$ such that $(p_{0},q_{0})\Vdash\operatorname{dom}(\sigma^{\prime})\subseteq\check{x}$. Now we can define a function $q_{1}$ on $x$ such that for $\gamma\in x$, $q_{1}(\gamma)$ is the $\operatorname{Add}(\tau,\gamma)$-name corresponding to $\sigma^{\prime}(\check{\gamma})$, which is an $\operatorname{Add}(\tau,\nu)$-name for an $\operatorname{Add}(\tau,\gamma\smallsetminus\nu)$-name. It follows that $(p_{0}\cup p_{1},q_{0}\cup q_{1})$ has the property that $\iota(p_{0}\cup p_{1},q_{0}\cup q_{1})\leq((p_{0},q_{0}),\sigma)$. ∎
@@
By a nice name argument, $\mathbb{M}_{0}(\tau,\mu,\kappa)$ forces $2^{\tau}=\kappa$. Because $\mathbb{M}_{0}(\tau,\mu,\kappa)$ is ${<}\,\tau$-directed closed, it preserves cardinals below and including $\tau$. Because we can project onto it from the product of a $\tau^{+}$-cc. and a ${<}\,\mu$-closed poset, it preserves cardinals in the interval $[\tau^{+},\mu]$. Lastly, every cardinal in $(\mu,\kappa)$ is collapsed to $\mu$ and cardinals above and including $\kappa$ are preserved by the $\kappa$-cc. ∎

### v2

@@ 4. Ineffability Witnesses
    $\mathbb{P}$ is of size $\leq\lambda$ and $\kappa$-$c$.,
@@
      $M\cap\lambda]^{<M\cap\kappa}\subseteq M$.
@@
      Whenever $G$ is $\mathbb{P}$-generic over $V$, $\pi[G\cap M]$ is $\pi(\mathbb{P})$-generic over $V$ and the pair $(V[\pi[G\cap M]],V[G])$ has the $\pi(\delta)-$approximation property.
@@
Without loss of generality, consider $\mathbb{P}$ to be a partial order on $\lambda$. Denote by $\langle\cdot,\cdot\rangle$ the Gödel pairing function.
Let $\dot{f}$ be a $\mathbb{P}$-name for a $\delta$-slender $(\kappa,\lambda)$-list, forced by some $p$. Let $\dot{F}$ be the function corresponding to the club (in some $H(\Theta^{\prime})]^{<\kappa}$) witnessing the $\delta$-slenderness of $\dot{f}$. Let $\Theta$ be large so that $H(\Theta)$ contains all relevant objects.
We will transform $\dot{f}$ into a ground-model $(\kappa,\lambda)$-list. To this end, let $a\in[\lambda]^{<\kappa}$. We consider two cases:
@@
    If there exists $M\prec H(\Theta)$ (with Mostowski-Collapse $\pi$) such that $M\cap\kappa\in\kappa$, $M\cap\lambda=a$, $M\cap\lambda]^{<M\cap\kappa}\subseteq M$, $\kappa,\lambda,\dot{f},\dot{F}\in M$ and for some $\pi(\mathbb{P})$-name $\dot{x}_{a}$ for a subset of $\pi[M\cap\lambda]$ and a condition $p_{a}\leq p$, $p_{a}\Vdash\dot{f}(M\cap\lambda)=\pi^{-1}[\dot{x}_{a}^{\pi[\Gamma\cap M]}]$, let
@@
Let $M\in[H(\Theta)]^{<\kappa}$ be a $\lambda$-ineffability witness for $\kappa$ with respect to $e$ as in the requirements. Let $\pi\colon M\to N$ be the Mostowski-Collapse and $a:=M\cap\lambda$ as well as $\theta:=\pi[a]$. We will show that case (1) holds. Let $G_{0}$ be $\mathbb{P}$-generic containing $p$ and $G_{0}^{\prime}:=\pi[G_{0}\cap M]$. By Lemma 4.6, $\pi$ extends to $\pi\colon M[G_{0}]\to N[G_{0}^{\prime}]$ and $\theta]^{<\nu}\cap V[G_{0}^{\prime}]\subseteq N[G_{0}^{\prime}]$. Furthermore, $M[G_{0}]\prec H(\Theta)^{V[G_{0}]}$, so $M[G_{0}]\cap H(\Theta^{\prime})$ is closed under $\dot{F}^{G_{0}}$ and therefore witnesses the slenderness of $\dot{f}^{G_{0}}$.
Assume $\pi[\dot{f}^{G_{0}}(a)]\notin V[G_{0}^{\prime}]$. By the $\pi(\delta)-$approximation property there is $z\in V[G_{0}^{\prime}]$ with ordertype $\pi(\delta)$ such that $\pi[\dot{f}^{G_{0}}(a)]\cap z\notin V[G_{0}^{\prime}]$. We can assume $z\subseteq\pi[a]$, so we have $z\in N[G_{0}^{\prime}]$ and $\pi^{-1}(z)=\pi^{-1}[z]\in M[G_{0}]$. We have
$\pi[\dot{f}^{G_{0}}(a)]\cap z=\pi[\dot{f}^{G_{0}}(a)]\cap\pi[\pi^{-1}[z]]=\pi[\dot{f}^{G_{0}}(a)\cap\pi^{-1}[z]]$
and, since $M[G_{0}]\cap H(\Theta^{\prime})$ witnesses the slenderness of $\dot{f}^{G_{0}}$, we have $\dot{f}^{G_{0}}(a)\cap\pi^{-1}[z]\in M[G_{0}]$, which implies that $\dot{f}^{G_{0}}(a)\cap\pi^{-1}[z]\subseteq M[G_{0}]$ and thus that
$\pi[\dot{f}^{G_{0}}(a)\cap\pi^{-1}[z]]=\pi(\dot{f}^{G_{0}}(a)\cap\pi^{-1}[z])\in N[G_{0}^{\prime}]$
which presents a contradiction. So $\pi[\dot{f}^{G_{0}}(a)]\in V[G_{0}^{\prime}]$. Ergo there is $p_{a}\leq p$ as well as a $\pi(\mathbb{P})$-name $\dot{x}_{a}$ for a subset of $\theta$ such that $p_{a}\Vdash\pi[\dot{f}(a)]=\dot{x}_{a}^{\pi[\Gamma\cap M]}$ which is what we wanted to show.
Now our aim is to show that $p_{a}$ forces $M[\Gamma]$ to be a $\lambda$-ineffability witness for $\kappa$ with respect to $\dot{f}$. So let $G_{1}$ be a $\mathbb{P}$-generic filter containing $p_{a}$. It is clear that $\kappa,\lambda,\dot{f}^{G_{1}}\in M[G_{1}]$. By assumption there is $b_{e}\in M$ such that $b_{e}\cap M=e(M\cap\lambda)$. Define
@@
all that is left is to show $b_{f}\cap M=\dot{f}^{G_{1}}(M\cap\lambda)=\dot{f}^{G_{1}}(M[G_{1}]\cap\lambda)$ (where the last equality holds by Lemma 4.6).
Let $\beta\in b_{f}\cap M$. By elementarity there is $\alpha\in G_{1}\cap M$ such that $\langle\alpha,\beta\rangle\in b_{e}$ and we have $\langle\alpha,\beta\rangle\in b_{e}\cap M=e(a)=\pi^{-1}[g(a)]$. So $\pi(\langle\alpha,\beta\rangle)=\langle\pi(\alpha),\pi(\beta)\rangle\in g(a)$. By the definition, $\pi(\alpha)\Vdash\pi(\check{\beta})\in\dot{x}_{a}$. Hence $\pi(\beta)\in\dot{x}_{a}^{\pi[G_{1}\cap M]}$ and $\beta\in\dot{f}^{G_{1}}(a)$.
Let $\beta\in\dot{f}^{G_{1}}(M\cap\lambda)$, so $\pi(\beta)\in\dot{x}_{a}^{\pi[G_{1}\cap M]}$. Thus there exists $\alpha^{\prime}\in\pi[G_{1}\cap M]$ such that $\alpha^{\prime}\Vdash\pi(\beta)\in\dot{x}_{a}$. Let $\alpha^{\prime}=\pi(\alpha)$, $\alpha\in G_{1}\cap M$. Hence $\pi(\langle\alpha,\beta\rangle)\in g(a)$ and $\langle\alpha,\beta\rangle\in e(a)=b_{e}\cap M$. Ergo $\beta\in b_{f}\cap M$. ∎
@@
Let $\mathbb{P}:=\operatorname{Add}(\tau,\kappa)$. Let $e$ be any $(\kappa,\lambda)$-list, $\Theta$ large and $x\in[H(\Theta)]^{<\kappa}$. Find a $\lambda$-ineffability witness $M$ for $\kappa$ as in Lemma 4.3 with $\tau^{+}+1\subseteq M$. It follows that $\pi(\mathbb{P})=\operatorname{Add}(\tau,M\cap\kappa)$. If $G$ is $\mathbb{P}$-generic, $\pi[G\cap M]$ is $\pi(\mathbb{P})$-generic and $V[G]$ is an extension of $V[\pi[G\cap M]]$ by $\operatorname{Add}(\tau,\kappa\smallsetminus(M\cap\kappa))$. In $V[\pi[G\cap M]]$, this forcing is $\tau^{+}$-Knaster and thus has the $\tau^{+}=\pi(\tau^{+})-$approximation property (viewing it as an ordering on the product $\mathbb{P}\times\{0\}$ which is clearly iteration-like). ∎
@@ 5. Controlling the Slenderness at $\kappa$
We will start with the easiest construction regarding $\operatorname{ISP}$: In this section we will define a forcing $\mathbb{M}_{0}(\tau,\mu,\kappa)$ which forces $2^{\tau}=\mu^{+}=\kappa$ and $\operatorname{ISP}(\tau^{+},\kappa,\lambda)$ (if $\kappa$ is $\lambda$-ineffable), answering Problem 4.31 given by Mohammadpour in [17]. Additionally, it will force $\neg\operatorname{ISP}(\tau,\kappa,\kappa)$. Thus we are able to control exactly how slender lists must be to have guaranteed ineffable branches.
For this section, fix regular cardinals $\tau<\mu<\kappa$ such that $\tau^{<\tau}=\tau$ and $\kappa$ is inaccessible.
For technical reasons, for any ordinal $\gamma$, we let $\operatorname{Add}(\tau,\gamma)$ be the $\tau-$supported product of $\operatorname{Add}(\tau)$ of length $\gamma$, i.e. conditions are partial functions $p$ on $\gamma$ of size $\tau$ such that for all $\alpha\in\operatorname{dom}(p)$, $p(\alpha)\in\operatorname{Add}(\tau)$, ordered in the usual way.
@@
    $q$ is a $\mu-$sized partial function on $\nu$ such that whenever $\gamma\in\operatorname{dom}(q)$, $\gamma=\delta+1$ for an inaccessible cardinal $\delta$ and $q(\gamma)$ is an $\operatorname{Add}(\tau,\gamma)$-name for a condition in $\dot{\operatorname{Coll}}(\check{\mu},\check{\delta})$.
@@
The forcing $\mathbb{M}_{0}$ was used by Levine in [16] to give an easier construction of a model with a disjoint stationary sequence on an arbitrary double successor cardinal and disjoint stationary sequences on two successive cardinals.
We can obviously view $\mathbb{M}_{0}(\tau,\mu,\nu)$ as an ordering on some product $\operatorname{Add}(\tau,\nu)\times\mathbb{T}(\tau,\mu,\nu)$, which is based.
@@
    $\mathbb{M}_{0}(\tau,\mu,\nu)$ is $\tau-$directed closed.
@@
    The term ordering on $\mathbb{M}_{0}(\tau,\mu,\nu)$ is $\mu-$closed,
@@
The statements (1) and (2) are standard.
We give the proofs for (3) and (4) in order to familiarize ourselves with the definitions.
Regarding the base ordering, we observe that $(p,q)\leq(p^{\prime},q)$ if and only if $p\leq p^{\prime}$ in the ordering on $\operatorname{Add}(\tau,\gamma)$. So the base ordering is equal to $\operatorname{Add}(\tau,\gamma)$ which is $(2^{<\tau})^{+}=\tau^{+}$-Knaster.
To show (4), let $(p,q_{\alpha})_{\alpha<\delta}$ (for $\delta<\mu$) be a descending sequence. Let $x:=\bigcup_{\alpha<\delta}\operatorname{dom}(q_{\alpha})$. $|x|<\mu$. On $x$, define a function $q$ as follows: Given $\gamma\in x$, let $\alpha_{0}$ be such that $\gamma\in\operatorname{dom}(q_{\alpha_{0}})$ (so $\gamma\in\operatorname{dom}(q_{\alpha})$ for all $\alpha\geq\alpha_{0})$. Then $(q_{\alpha}(\gamma))_{\alpha_{0}\leq\alpha<\delta}$ is forced by $p$ to be a descending sequence in $\dot{\operatorname{Coll}}(\check{\mu},\check{\delta})$, so by the maximal principle we can fix $q(\gamma)$ which is forced to be a lower bound. It follows that $(p,q)$ is a lower bound of $(p,q_{\alpha})_{\alpha<\delta}$.
Regarding the projection property, given $(p^{\prime},q^{\prime})\leq(p,q)$, let $q^{\prime\prime}$ be a function with domain $\operatorname{dom}(q^{\prime})$ such that for all $\alpha\in\operatorname{dom}(q^{\prime})$, $q^{\prime\prime}(\alpha)$ is forced by $p^{\prime}$ to be equal to $q^{\prime}(\alpha)$ and by conditions incompatible to $p^{\prime}$ to be equal to $q(\alpha)$ (or $\emptyset$, if $q(\alpha)$ is not defined). Clearly, $q^{\prime\prime}$ is as required.
$\mathbb{M}_{0}$ has the refinement property since stronger conditions force more.
For the mixing property, let $(p,q_{0}),(p,q_{1})\leq(p,q)$. Let $p_{0},p_{1}$ be extensions of $p$ such that $p_{0}(0)$ and $p_{0}(1)$ are incompatible. Construct a function $q^{\prime}$ with domain $x:=\operatorname{dom}(q_{0})\cup\operatorname{dom}(q_{1})$ as follows: Let $\alpha\in x$. Then $p_{0}\upharpoonright\alpha$ and $p_{1}\upharpoonright\alpha$ are incompatible since $\alpha$ is the successor of a cardinal. So we can let $q^{\prime}(\alpha)$ be such that it is forced by $p_{0}\upharpoonright\alpha$ to be equal to $q_{0}(\alpha)$ (or $\emptyset$, if $q_{0}(\alpha)$ is not defined) and by conditions incompatible with $p_{0}\upharpoonright\alpha$ (in particular, $p_{1}\upharpoonright\alpha$) to be equal to $q_{1}(\alpha)$ (or $\emptyset$, if $q_{1}(\alpha)$ is not defined). It follows that $(p,q^{\prime})\leq(p,q)$ and for $i=0,1$, $(p_{i},q^{\prime})\leq(p,q_{i})$. ∎
@@
    $q$ is a $\mu-$sized partial function on $\kappa\smallsetminus\nu$ such that for all $\gamma\in\operatorname{dom}(q)$, $\gamma=\delta+1$ for an inaccessible cardinal $\nu$ and $q(\gamma)$ is an $\operatorname{Add}(\tau,\gamma\smallsetminus\nu)$-name for a condition in $\dot{\operatorname{Coll}}(\check{\mu},\check{\nu})$.
@@
    $\mathbb{M}_{0}(G,\tau,\mu,\kappa\smallsetminus\nu)$ is $\tau-$directed closed.
@@
    The term ordering on $\mathbb{M}_{0}(G,\tau,\mu,\kappa\smallsetminus\nu)$ is $\mu-$closed,
@@
Let $\nu\in(\mu,\kappa)$ be inaccessible and let $G$ be $\mathbb{M}_{0}(\tau,\mu,\nu)$-generic. In $V[G]$, the forcing $\mathbb{M}_{0}(G,\tau,\mu,\kappa\smallsetminus\nu)$ has the $\tau^{+}-$approximation property.
@@
As is common, we define $\iota((p,q)):=((p\upharpoonright\nu,q\upharpoonright\nu),(\check{p}\upharpoonright[\nu,\kappa),\overline{q}))$, where $\overline{q}$ is $\mathbb{M}_{0}(\tau,\mu,\nu)$-name for a function on $\kappa\smallsetminus\nu$ given by recomputing $q\upharpoonright[\nu,\kappa)$.
It is clear that $\iota$ preserves $\leq$. Let $((p_{0},q_{0}),\sigma)$ be a condition in $\mathbb{M}_{0}(\tau,\mu,\nu)*\mathbb{M}_{0}(\Gamma,\tau,\mu,\kappa\smallsetminus\nu)$. By strengthening $(p_{0},q_{0})$ if necessary, we can assume that it forces $\sigma=(\check{p_{1}},\sigma^{\prime})$ for some $p_{1}\in\operatorname{Add}(\tau,\kappa\smallsetminus\nu)$ (because $\mathbb{M}_{0}(\tau,\mu,\nu)$ is $\tau-$directed closed). $\sigma^{\prime}$ is forced to be a $\mu-$sized sequence of $\mu-$sized subsets of $V$, so we can (again, strengthening $(p_{0},q_{0})$ if necessary) assume that $\sigma^{\prime}$ is an $\operatorname{Add}(\tau,\nu)$-name. Lastly, by the $\mu$-cc. of $\operatorname{Add}(\tau,\nu)$, there is a set $x\in V$ such that $(p_{0},q_{0})\Vdash\operatorname{dom}(\sigma^{\prime})\subseteq\check{x}$. Now we can define a function $q_{1}$ on $x$ such that for $\gamma\in x$, $q_{1}(\gamma)$ is the $\operatorname{Add}(\tau,\gamma)$-name corresponding to $\sigma^{\prime}(\check{\gamma})$, which is an $\operatorname{Add}(\tau,\nu)$-name for an $\operatorname{Add}(\tau,\gamma\smallsetminus\nu)$-name.
It follows that $(p_{0}\cup p_{1},q_{0}\cup q_{1})$ has the property that $\iota(p_{0}\cup p_{1},q_{0}\cup q_{1})\leq((p_{0},q_{0}),\sigma)$. ∎
@@
By a nice name argument, $\mathbb{M}_{0}(\tau,\mu,\kappa)$ forces $2^{\tau}=\kappa$: Any $\mathbb{M}_{0}(\tau,\mu,\kappa)$-name for a subset of $\tau$ be can viewed as function from $\tau$ into the set of maximal antichains in $\mathbb{M}_{0}(\tau,\mu,\kappa)$. Since $\mathbb{M}_{0}(\tau,\mu,\kappa)$ is $\kappa$-cc and has size $\kappa$, that set has size $\kappa^{<\kappa}=\kappa$ by the inaccessibility of $\kappa$. By the same argument, there are at most $\kappa$ functions from $\tau$ into the set of maximal antichains in $\mathbb{M}_{0}(\tau,\mu,\kappa)$ and so there are (up to equivalence) at most $\kappa$ many $\mathbb{M}_{0}(\tau,\mu,\kappa)$-names for subsets of $\tau$. However, $\mathbb{M}_{0}(\tau,\mu,\kappa)$ clearly projects onto $\operatorname{Add}(\tau,\kappa)$ and therefore forces $2^{\tau}\geq\kappa$.
Because $\mathbb{M}_{0}(\tau,\mu,\kappa)$ is $\tau-$directed closed, it preserves cardinals below and including $\tau$. Because we can project onto it from the product of a $\tau^{+}$-cc. and a $\mu-$closed poset, it preserves cardinals in the interval $\tau^{+},\mu]$. Lastly, every cardinal in $(\mu,\kappa)$ is collapsed to $\mu$ and cardinals above and including $\kappa$ are preserved by the $\kappa$-cc. ∎